Complete the square. x^2= -12 - 10x
Answer by stanbon(75887)

Complete the square. x^2= -12 - 10x
-------------
x^2+10x + ? = -12
---
x^2+10x + 5^2 = -12+25
----
(x+5)^2 = 13
===============
Cheers,
Stan H.